OpenStudy (anonymous):

−10t5 + 15t4 + 9t3 factor

OpenStudy (theviper):

Here \[t^3 \]is common.

OpenStudy (anonymous):

Oh, we've done these quadratic forms before . . . ;-)

OpenStudy (theviper):

\[\LARGE{\color{red}{t^3(-10t^2+15t+9)}}\] Then factor the eq in the brackets:)

OpenStudy (anonymous):

\[−10t^5 + 15t^4 + 9t^3\] \[t^3(−10t^2 + 15t + 9)\] now just use the quadratic formula

OpenStudy (anonymous):

\[(-b +/- \sqrt{(b^2 -4ac))})/2a\]

OpenStudy (anonymous):

Quadratic formula isn't required for factoring. I think the quadratic part is actually prime.

OpenStudy (theviper):

\[\Huge{\color{blue}{\frac{-b- \sqrt{b^2-4ac}}{2a}}}\]

OpenStudy (theviper):

I think now u can do it, can't u?:)

OpenStudy (theviper):

\[\Huge{\color{orange}{a=-10,b=15,c=9}}\]

OpenStudy (anonymous):

well now im confused..lol

OpenStudy (anonymous):

so itd be( -15 - sq root of 15^2 -4(10*9) )/2(10)

OpenStudy (theviper):

First we took \[t^3\]as common. Then only u have to factorise the eq in the brackets by the quadratic formula I & @saljudieh07 have given:)

OpenStudy (theviper):

Clear now??

OpenStudy (anonymous):

*Using the quadratic formula -is not- factoring. Using the quadratic formula is wrong here.

OpenStudy (anonymous):

yeh i think cliff is right, because the quadratic formula would find the values of t, if the quadratic equation is equal to zero. so in this case the final answer is: \[t^3(−10t2+15t+9)\]

OpenStudy (anonymous):

Taking a*c = (-10)(9) = -90; there are no integer factors of -90 that add or subtract to make b=15, so -10t^2+15t+9 is prime.

OpenStudy (anonymous):

ok..ill go with that

OpenStudy (ledah):

Even...this confused me.. >.>